# PKS 2131--021 – Discovery of Strong Coherent Sinusoidal Variations from Radio to Optical Frequencies: Compelling Evidence for a Blazar Supermassive Black Hole Binary

### v1

@@ ACKNOWLEDGEMENTS
We thank the California Institute of Technology and the Max Planck Institute for Radio Astronomy for supporting the OVRO 40 m program under extremely difficult circumstances over the last 10 years in the absence of agency funding. Without this private support these observations could not have been made. We also thank all the volunteers who have enabled this work to be carried out. Prior to 2016, the OVRO program was supported by NASA grants NNG06GG1G, NNX08AW31G, NNX11A043G, and NNX13AQ89G from 2006 to 2016 and NSF grants AST-0808050 and AST-1109911 from 2008 to 2014. S.K. & K.T. acknowledge support from the European Research Council (ERC) under the European Unions Horizon 2020 research and innovation programme under grant agreement No. 771282. I.L and S.K. were funded by the European Union ERC-2022-STG - BOOTES - 101076343. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. W.M. acknowledges support from ANID project Basal FB210003. A.S. and R.B acknowledge support by a grant from the Simons Foundation (00001470,RB,AS). R.R. acknowledges support from ANID Basal FB210003 and ANID-FONDECYT grant 1181620. C.O. acknowledges support from the Natural Sciences and Engineering Research Council (NSERC) of Canada. T.H. acknowledges support from the Academy of Finland projects 317383, 320085, 322535, and 345899. The NANOGrav collaboration receives support from National Science Foundation (NSF) Physics Frontiers Center award numbers 1430284 and 2020265. Part of this research was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2011.0.00001.CAL. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology.
Appendix A KO model Parameters for the Sinusoidally Varying Core Component
@@
Homan et al. (2021) were specifically interested in the PKS 2131-021 component brightness temperatures and they fitted all of the core regions in a consistent fashion, taking particular care not to mix components between epochs. The details of their approach are given in the methods section of their paper. They found a median $T_{\rm b,\,observed}=5.8\times 10^{11}$ over 24 epochs from 1995 to 2013. They also found an estimated intrinsic median brightness temperature for their sample of $T_{\rm b,\,intrinsic}=4.1\times 10^{10}$, which gives an estimated Doppler factor of $D=T_{\rm b,\,observed}/T_{\rm b,\,intrinsic}=14$. Thus, using the value $\theta=3.8^{\circ}$ determined by Homan et al. (2021), for this paper we adopt the value $\beta=0.995$, corresponding to $\gamma=10$ and $D=14$. The amplitude of the 15 GHz OVRO light curve, assuming a fixed period $P=1739.2$ days for the sinusoidal component, is 0.42 Jy. In order not to detect it, we need to suppress the amplitude of the first harmonic to below 0.1×0.1$\times$0.1 × the amplitude of the fundamental. Hence the amplitude of the first harmonic should be less than 42 mJy. A sinusoidal first harmonic of amplitude 42 mJy is shown in Fig. 16. To achieve this suppression of the amplitude of the first harmonic, we need a value of $\beta_{1}\ll 1$, so we adopt a value $\beta_{1}=0.020$.
@@
$\psi_{{}_{\rm eq}}=1.74\,r^{-{1\over 17}}S^{8\over 17}\nu^{-{{35+2\alpha}\over 34}}(1+z)^{{15-2\alpha}\over 34}F(\alpha)\,,$ (A1)
@@
$T_{\rm eq}=5.84\times 10^{11}\biggl{[}{{r\over{(1+z)}}}\biggr{]}^{2/17}F(\alpha)^{-2}(1+z)^{(2\alpha-13)/17}S^{1/17}\bigl{(}10^{3}\nu\bigr{)}^{(1+2\alpha)/17}{\rm K}\,.$ (A2)
@@
Appendix B The Absence of Higher Harmonics
@@
$\ln\mathcal{L}=-\frac{1}{2}\sum_{i}^{t<45500}\left(\frac{(S_{i}-S(t_{i}))^{2}}{\sigma_{i}^{2}+\sigma_{1}^{2}}+\ln(\sigma_{i}^{2}+\sigma_{1}^{2})\right)-\frac{1}{2}\sum_{i}^{t>52800}\left(\frac{(S_{i}-S(t_{i}))^{2}}{\sigma_{i}^{2}+\sigma_{2}^{2}}+\ln(\sigma_{i}^{2}+\sigma_{2}^{2})\right).$ (B1)
@@
Figure 16: A first harmonic of amplitude 0.042 Jy is shown by the red curve. Sinusoidal first harmonic fluctuations of this amplitude cannot be ruled out by our data (see Appendix A), because of the non-Gaussian nature of the noise, which clearly has long-term correlated flucutations at this level . Other symbols and curves are the same as those in Fig. 13.
@@
In order to test for a possible first harmonic, we carried out the following exercise. We added an artificial sinusoidal signal with a period of $P/2$, and an amplitude of 0.012 Jy to the data. We then fitted sine-wave models (with and without the first harmonic) to the synthetic data. The best-fit amplitude of the first harmonic was $0.025\pm 0.006$ Jy, and the BIC of the model with the first harmonic was ${\rm BIC}=-2661.7$. The BIC for the model without harmonics was $-2665.0$, so $\delta({\rm BIC})=+3.3$. When we added an artificial sinusoidal signal with an amplitude of 0.025 Jy, the best-fit amplitude of the first harmonic was $0.036\pm 0.006$ Jy, and $\delta({\rm BIC})=-31.3$. In conclusion, we find that an amplitude of the first harmonic of 0.025 Jy is just consistent (at the $2\sigma$ level) with the data assuming random variations. Given that the variations are clearly not random, as can be seen by the black points in Fig. 16, which show long-term correlated fluctuations, the level of consistency is considerably better than this. So a first harmonic of amplitude 42 mJy cannot be ruled out by the data.
Appendix C Sine-wave fits to the optical data
We fitted the following sine-wave model to the optical data:
$m(t)=A\sin(\phi-\phi_{0})+m_{0},$ (C1)
$\ln\mathcal{L}=-\frac{1}{2}\sum_{g,r,i}\sum_{j=1}^{N_{\rm obs}}\left(\frac{(m_{j}-m_{j}^{\rm model})^{2}}{\sigma_{j}^{2}+\sigma_{0}^{2}}+\ln(\sigma_{j}^{2}+\sigma_{0}^{2})\right).$ (C2)
Table 5: Results of the sine-wave fits to the optical data
Parameter Value
$P$ (d) $1764\pm 36$ $1762\pm 36$
$\phi_{0}$ (rad) $4.372\pm 0.056$
$\phi_{0}$ ($r$) $4.353\pm 0.080$
$\phi_{0}$ ($g$) $4.349\pm 0.084$
$\phi_{0}$ ($i$) $4.456\pm 0.113$
$m_{0}$ ($r$) $17.454\pm 0.021$ $17.453\pm 0.021$
$A$ ($r$) $0.358\pm 0.032$ $0.358\pm 0.032$
$\sigma_{0}$ ($r$) $0.363\pm 0.015$ $0.363\pm 0.015$
$m_{0}$ ($g$) $18.030\pm 0.022$ $18.030\pm 0.022$
$A$ ($g$) $0.366\pm 0.033$ $0.366\pm 0.034$
$\sigma_{0}$ ($g$) $0.383\pm 0.016$ $0.383\pm 0.016$
$m_{0}$ ($i$) $17.146\pm 0.035$ $17.155\pm 0.037$
$A$ ($i$) $0.379\pm 0.055$ $0.388\pm 0.058$
$\sigma_{0}$ ($i$) $0.376\pm 0.023$ $0.376\pm 0.023$
$\chi^{2}/\mathrm{d.o.f.}$ $785.4/780$ $786.7/778$

### v2

@@ ACKNOWLEDGEMENTS
This work is supported by NSF grants AST2407603 and AST2407604. We thank the California Institute of Technology and the Max Planck Institute for Radio Astronomy for supporting the OVRO 40 m program under extremely difficult circumstances over the last 8 years in the absence of agency funding. Without this private support these observations could not have been made. We thank Adam Hincks for useful discussions. We also thank all the volunteers who have enabled this work to be carried out. Prior to 2016, the OVRO program was supported by NASA grants NNG06GG1G, NNX08AW31G, NNX11A043G, and NNX13AQ89G from 2006 to 2016 and NSF grants AST-0808050 and AST-1109911 from 2008 to 2014. S.K. & K.T. acknowledge support from the European Research Council (ERC) under the European Unions Horizon 2020 research and innovation programme under grant agreement No. 771282. I.L and S.K. were funded by the European Union ERC-2022-STG - BOOTES - 101076343. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. W.M. acknowledges support from ANID project Basal FB210003. A.S. and R.B acknowledge support by a grant from the Simons Foundation (00001470,RB,AS). R.R. and B.M. and P.V.d.l.P. acknowledge support from ANID Basal AFB-170002, from Núcleo Milenio TITANs (NCN2023_002), and CATA BASAL FB210003. P.V.d.l.P. also acknowledges support by the National Agency for Research and Development (ANID) / Scholarship Program / Doctorado Nacional/2023–21232103. C.O. acknowledges support from the Natural Sciences and Engineering Research Council (NSERC) of Canada. T.H. acknowledges support from the Academy of Finland projects 317383, 320085, 322535, and 345899. The NANOGrav collaboration receives support from National Science Foundation (NSF) Physics Frontiers Center award numbers 1430284 and 2020265. Part of this research was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2011.0.00001.CAL. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology.
Appendix A Sine-wave fits to the optical data
We fitted the following sine-wave model to the optical data:
$m(t)=A\sin(\phi-\phi_{0})+m_{0},$ (A1)
$\ln\mathcal{L}=-\frac{1}{2}\sum_{g,r,i}\sum_{j=1}^{N_{\rm obs}}\left(\frac{(m_{j}-m_{j}^{\rm model})^{2}}{\sigma_{j}^{2}+\sigma_{0}^{2}}+\ln(\sigma_{j}^{2}+\sigma_{0}^{2})\right).$ (A2)
Table 4: Results of the sine-wave fits to the optical data
Parameter Value
$P$ (d) $1764\pm 36$ $1762\pm 36$
$\phi_{0}$ (rad) $4.372\pm 0.056$
$\phi_{0}$ ($r$) $4.353\pm 0.080$
$\phi_{0}$ ($g$) $4.349\pm 0.084$
$\phi_{0}$ ($i$) $4.456\pm 0.113$
$m_{0}$ ($r$) $17.454\pm 0.021$ $17.453\pm 0.021$
$A$ ($r$) $0.358\pm 0.032$ $0.358\pm 0.032$
$\sigma_{0}$ ($r$) $0.363\pm 0.015$ $0.363\pm 0.015$
$m_{0}$ ($g$) $18.030\pm 0.022$ $18.030\pm 0.022$
$A$ ($g$) $0.366\pm 0.033$ $0.366\pm 0.034$
$\sigma_{0}$ ($g$) $0.383\pm 0.016$ $0.383\pm 0.016$
$m_{0}$ ($i$) $17.146\pm 0.035$ $17.155\pm 0.037$
$A$ ($i$) $0.379\pm 0.055$ $0.388\pm 0.058$
$\sigma_{0}$ ($i$) $0.376\pm 0.023$ $0.376\pm 0.023$
$\chi^{2}/\mathrm{d.o.f.}$ $785.4/780$ $786.7/778$
Appendix B Fit of sinusoidal model using simulation-based likelihood
As shown in Fig. 2, the OVRO and Haystack 15 GHz lightcurves appear to fit the same sine wave, albeit with different amplitudes and zero points. We can model these as
$S_{i}^{O}$ $=A^{O}\sin(2\pi[(t_{i}^{O}-t_{0})/P]+\phi)+S_{0}^{O}+n_{i}^{O}$
$S_{i}^{H}$ $=A^{H}\sin(2\pi[(t_{i}^{H}-t_{0})/P]+\phi)+S_{0}^{H}+n_{i}^{H}$ (B1)
Where $i$ is a sample index (different range for OVRO and Haystack), $t_{i}$ is is the observation time of sample $i$, and $n_{i}$ its noise, and the superscripts $O$ and $H$ refer to OVRO and Haystack respectively. The unknown parameters are the period $P$, phase $\phi$, amplitudes $A^{O}$ and $A^{H}$ and the flux density offsets $S_{0}^{O}$ and $S_{0}^{H}$.
B.1 The noise covariance matrix
Figure 16: Correlation part of the sample-sample covariance matrix $N^{O}$ for the OVRO lightcurve. Dark blue is zero, dark red is one. The blocky structure is caused by the uneven sample spacing.
Figure 17: $\log P$ plotted as a function of $A^{HO}$ (horizontal axis, the Haystack amplitude relative to OVRO) and $P$ (vertical axis, the sine period in days). The black contours are steps of 1 in $\log P$.
The noise includes the instrument and atmospheric noise and any AGN variability that is not part of the sinusoidal variation. We approximate the noise as Gaussian, with covariances $N^{O}$ and $N^{H}$. These covariance matrices are challenging to estimate due to the limited amount of data available, but we can use as a proxy the lightcurve simulations. Hence we build the sample-sample covariance matrixes $N^{O}_{ij}$ and $N^{H}_{ij}$. These covariance matrices originally had some unphysical properties caused by limitations in the simulation methodology: As the point separation grows, the correlation falls as expected, but eventually becomes negative before starting to rise again. The rise is caused by the periodic boundary conditions implied by the Fourier methods used when constructing the simulations, while we interpret the negative correlation region as being caused by the characteristic length scale of the sinusoidal signal itself. Neither of these properties is relevant to the AGN red noise we wish to model, so we modified the covariance matrices to set the negative and rising areas equal to zero. [^3] We have confirmed that the unphysical periodicity of the simulations did not impact the significance estimate the simulations were originally built for. The resulting OVRO covariance matrix is shown in Fig. 17.
B.2 Likelihood
The model in equation B1 is inefficient to fit because it has four non-linear parameters: $P$, $\phi$, $A^{O}$ and $A^{H}$. We can reduce this to two by re-parameterizing as follows:
$S_{i}^{O}$ $=\alpha\sin(\theta_{i})+\beta\cos(\theta_{i})+S_{0}^{O}+n_{i}^{O}$
$S_{i}^{H}$ $=A^{HO}\Big{(}\alpha\sin(\theta_{i})+\beta\cos(\theta_{i})\Big{)}+S_{0}^{H}+n_{i}^{H}$ (B2)
where $\theta_{i}=2\pi(t_{i}-t_{0})/P$, and $A^{HO}=A^{H}/A^{O}$ is the ratio of the Haystack amplitude to the OVRO amplitude. With this parameterization the only non-linear parameters are $P$ and $A^{HO}$. Given these parameters, the rest can be evaluated in a single, linear step. We can write the model in linear algebra form
$S$ $=Qa+n$ (B3)
Here $S$ and $n$ are stacks of the OVRO and Haystack data and noise respectively; $a=[\alpha,\beta,S_{0}^{O},S_{0}^{H}]$ are the set of linear parameters; and $Q$ is a matrix that depends on $P$ and $A^{HO}$ and implements equations B3 and B4. We can then write the Likelihood as
$-2\log L$ $=|2\pi N|+(S-Qa)^{T}N^{-1}(S-Qa)$
$=|2\pi N|+(a-\hat{a})^{T}A^{-1}(a-\hat{a})+S^{T}N^{-1}S-\hat{a}A^{-1}\hat{a}$ (B4)
where $\hat{a}=(Q^{T}N^{-1}Q)^{-1}Q^{T}N^{-1}S$ and $\hat{A}^{-1}=Q^{T}N^{-1}Q$. These parameters have simple interpretations: $\hat{a}$ is the maximum-likelihood estimator for the linear parameters, and $\hat{A}$ is the covariance of this estimator. To reduce the non-linear fit to just two parameters we must marginalize over the linear parameters, which means integrating over all values of $a$. Since $a$ has a Gaussian likelihood with covariance $\hat{A}$, we know that its integral is $\sqrt{|2\pi\hat{A}|}$, so the marginalized likelihood is
$-2\log L^{\prime}$ $=|2\pi N|-|2\pi\hat{A}|+S^{T}N^{-1}S-\hat{a}^{T}\hat{A}^{-1}\hat{a}$ (B5)
The first and third terms don’t depend on any of our parameters and so can be ignored in an MCMC fit. Additionally ignoring the second term is equivalent to a Jeffrey’s minimum-information prior on $a$. We are therefore left with the following posterior probability for the non-linear parameters.
$\log P$ $=\frac{1}{2}\hat{a}^{T}\hat{A}^{-1}\hat{a}+\log\textrm{prior}$ (B6)
where “prior” is a manual prior on the parameters. We used $P\in[1000,3000]$ and $A^{HO}\in[0.1,4]$ in our fit.
The $P,A^{HO}$ posterior is shown in Fig. 17.
B.3 Full MCMC search
We performed a full MCMC search of the posterior in equation B6 using emcee with 4 walkers, a burn-in of 30 000 samples and 30 000 output samples. For each non-linear parameter sampled this way, we drew a sample of the linear parameters from their conditional distribution. Fig. 18 shows a corner plot for the resulting samples. Our final fit parameters are $P=1739.8\pm 17.4$ days, $\phi=1.078\pm 0.025$ radians offset from MJD $t_{0}=51000$, $A^{O}=0.374\pm 0.083$ Jy, $A^{H}=0.853\pm 0.304$.
Figure 18: Distribution of samples from MCMC fit of the parameters in equation B1. Note: Unlike elsewhere in the text, $\phi$ is here reported in units of periods instead of radians, so it’s smaller by a factor $2\pi$.
This fit is quite sensitive to the details of the noise covariance matrix. This matrix could be improved by re-estimating the noise properties after subtracting the best-fit signal model. As it is, the noise matrix is somewhat contaminated by the sinusoidal signal. We leave this for a future improvement, and for now note that the error bars are likely to be overestimated.
Appendix C KO model Parameters for the Sinusoidally Varying Core Component
@@
Homan et al. (2021) were specifically interested in the PKS 2131--021 component brightness temperatures and they fitted all of the core regions in a consistent fashion, taking particular care not to mix components between epochs. The details of their approach are given in the methods section of their paper. They found a median $T_{\rm b,\,observed}=5.8\times 10^{11}$ over 24 epochs from 1995 to 2013. They also found an estimated intrinsic median brightness temperature for their sample of $T_{\rm b,\,intrinsic}=4.1\times 10^{10}$, which gives an estimated Doppler factor of $D=T_{\rm b,\,observed}/T_{\rm b,\,intrinsic}=14$. Thus, using the value $\theta=3.8^{\circ}$ determined by Homan et al. (2021), for this paper we adopt the value $\beta=0.995$, corresponding to $\gamma=10$ and $D=14$. The amplitude of the 15 GHz OVRO light curve, assuming a fixed period $P=1739.2$ days for the sinusoidal component, is 0.42 Jy. In order not to detect it, we need to suppress the amplitude of the first harmonic to below 0.1×0.1$\times$0.1 × the amplitude of the fundamental. Hence the amplitude of the first harmonic should be less than 42 mJy. A sinusoidal first harmonic of amplitude 42 mJy is shown in Fig. 20. To achieve this suppression of the amplitude of the first harmonic, we need a value of $\beta_{1}\ll 1$, so we adopt a value $\beta_{1}=0.020$.
@@
$\psi_{{}_{\rm eq}}=1.74\,r^{-{1\over 17}}S^{8\over 17}\nu^{-{{35+2\alpha}\over 34}}(1+z)^{{15-2\alpha}\over 34}F(\alpha)\,,$ (C1)
@@
$T_{\rm eq}=5.84\times 10^{11}\biggl{[}{{r\over{(1+z)}}}\biggr{]}^{2/17}F(\alpha)^{-2}(1+z)^{(2\alpha-13)/17}S^{1/17}\bigl{(}10^{3}\nu\bigr{)}^{(1+2\alpha)/17}{\rm K}\,.$ (C2)
@@
Appendix D The Absence of Higher Harmonics
@@
$\ln\mathcal{L}=-\frac{1}{2}\sum_{i}^{t<45500}\left(\frac{(S_{i}-S(t_{i}))^{2}}{\sigma_{i}^{2}+\sigma_{1}^{2}}+\ln(\sigma_{i}^{2}+\sigma_{1}^{2})\right)-\frac{1}{2}\sum_{i}^{t>52800}\left(\frac{(S_{i}-S(t_{i}))^{2}}{\sigma_{i}^{2}+\sigma_{2}^{2}}+\ln(\sigma_{i}^{2}+\sigma_{2}^{2})\right).$ (D1)
@@
Figure 20: A first harmonic of amplitude 0.042 Jy is shown by the red curve. Sinusoidal first harmonic fluctuations of this amplitude cannot be ruled out by our data (see Appendix C), because of the non-Gaussian nature of the noise, which clearly has long-term correlated fluctuations at this level . Other symbols and curves are the same as those in Fig. 13.
@@
In order to test for a possible first harmonic, we carried out the following exercise. We added an artificial sinusoidal signal with a period of $P/2$, and an amplitude of 0.012 Jy to the data. We then fitted sine-wave models (with and without the first harmonic) to the synthetic data. The best-fit amplitude of the first harmonic was $0.025\pm 0.006$ Jy, and the BIC of the model with the first harmonic was ${\rm BIC}=-2661.7$. The BIC for the model without harmonics was $-2665.0$, so $\delta({\rm BIC})=+3.3$. When we added an artificial sinusoidal signal with an amplitude of 0.025 Jy, the best-fit amplitude of the first harmonic was $0.036\pm 0.006$ Jy, and $\delta({\rm BIC})=-31.3$. In conclusion, we find that an amplitude of the first harmonic of 0.025 Jy is just consistent (at the $2\sigma$ level) with the data assuming random variations. Given that the variations are clearly not random, as can be seen by the black points in Fig. 20, which show long-term correlated fluctuations, the level of consistency is considerably better than this. So a first harmonic of amplitude 42 mJy, i.e. one tenth of the fundamental, cannot be ruled out by the data.